The point (–4, 5) is the vertex of a square and one of its diagonals is 7x – y + 8 = 0. The equation of the other diagonal is
(A) 7x – y + 23 = 0 (B) 7y + x = 30 (C) 7y + x = 31 (D) x – 7y = 30

Grade:

1 Answers

Vikas TU
3 years ago
Equation of perpendicular line to
 7x−y+8=0
7x−y+8=0 
is x+7y=hx+7y=h 
which is passes through 
(−4, 5)(−4, 5) 
h = 31